# Temporal Analysis and Repair of Flaky Dockerfiles

### v1

@@ Abstract
Dockerfile flakiness, characterized by inconsistent build behavior without Dockerfile or project source code changes, poses significant challenges in Continuous Integration and Delivery (CI/CD) pipelines. This issue can lead to unreliable deployments and increased debugging efforts, yet it remains underexplored in current research. We conduct a systematic analysis of Dockerfile flakiness, presenting a comprehensive taxonomy of common flakiness categories, including dependency-related errors and server connectivity issues. Furthermore, we introduce FlakiDock, a tool leveraging large language models and retrieval-augmented generation techniques with dynamic analysis and an iterative feedback loop to automatically repair flaky Dockerfiles. Our evaluation shows that FlakiDock achieves a 73.55% repair accuracy, outperforming existing tools such as PARFUM by 12,581% and GPT-4-based prompting by 94.63%. These results underscore the effectiveness of FlakiDock in addressing Dockerfile flakiness and improving build reliability.
@@
Large Language Models, Few-shot Learning, Docker, Flakiness, Automated Program Repair
@@ I Introduction
Docker [^1] https://www.docker.com is a set of platform-as-a-service (PaaS) products that utilize OS-level virtualization to automate the building, deployment, and delivery of applications in environments known as containers. Docker images, created based on instructions specified in a Dockerfile, act as the blueprints for these containers. These images include all necessary components to execute an application, such as the code, runtime environment, libraries, and system tools. This approach allows developers to package their applications once and deploy them across various systems, streamlining the software deployment process. Dockerfiles play a pivotal role in the Continuous Integration and Continuous Delivery (CI/CD) pipeline, making it essential to ensure their reliability during the build stage to maintain software dependability.
Given the extensive reliance of Dockerfiles on various dependencies, packages, and configurations, ensuring their reliability becomes a critical concern. Existing academic [1, 2, 3, 4, 5, 6, 7, 8] and industrial efforts mainly focus on identifying Dockerfile smell or bug patterns. Dockerfile smells refer to an indication of potential issues or violations of best practices within a Dockerfile. Static analysis studies such as [2, 9, 10, 7] offer a linter to detect smells inside Dockerfiles. While these tools pinpoint a variety of Dockerfile smells, they rely on a set of pre-defined rules, often not well-maintained [10], to detect issues. A more recent tool called Parfum [4] takes a step further and enriches Dockerfile AST by incorporating structural information from the command lines to automatically detect and repair smells.
Despite the extensive investigations on Dockerfile smells, a notable gap exists in the current body of research: flakiness in Dockerfiles has not been comprehensively explored. The term flakiness has been widely used in research and industry to refer to tests with inconsistent results, alternating between passing and failing without any changes to the underlying code or environment. A flaky test can be considered as a bug in the test code, prone to generating unpredictable outcomes [11]. While flakiness is commonly used to illustrate non-deterministic behaviors in software testing, its scope can extend beyond this context. Similar to test flakiness [11], we define Dockerfile flakiness as the unpredictable behavior of the Dockerfile build process, where builds may succeed or fail overtime under identical Dockerfile context.
@@
Existing Dockerfile-related studies [1, 2, 3, 4, 5, 6, 7, 9, 10] assume the reliability of the Dockerfiles and their build output over time, so that a build failure indicates the changes applied directly to the source Dockerfile. Flakiness in Dockerfiles can cause several problems during the build and deployment of the projects such as disrupting CI/CD cycles by impeding automatic builds which causes delays or wasting time and effort of developers working on pinpointing and rectifying the cause of Dockerfile build flakiness.
@@
In this work, we examine the consistency and reliability of Dockerfile builds over time. In a longitudinal study, which spanned over a nine-month timeframe, we analyzed the build of 8,132 Dockerfiles and observed that 798 (9.81%) exhibits flakiness. Based on the observed instances of flakiness, we developed a taxonomy for characterizing Dockerfile flakiness. Our findings indicate that the most common types of flakiness are related to dependencies, web server connectivity, and security/authentication issues. Furthermore, we found that existing techniques, such as Parfum [4] an automated tool for repairing Dockerfile smells, can repair less than 1% of the occurrences of flakiness. To overcome the limitations of current tools, we propose a technique called FlakiDock to automatically repair Dockerfile flakiness.
@@
  • We conduct the first study to characterize Dockerfile flakiness. We develop the first taxonomy of Dockerfile flakiness by building Dockefiles and analyzing build outputs of 8,132 Dockerized projects.
  • We create a dataset called Flake4Dock including 798 and 7,325 flaky and non-flaky Dockerfiles. Flaky Dockerfiles are accompanied by categorization and build errors. We also provide repair information for 100 flaky Dockerfiles, specifically designed to evaluate flakiness detection and repair tasks.
  • We present a technique called FlakiDock that leverages both static and dynamic information from Dockerfiles to repair flakiness.
  • We perform an empirical evaluation of FlakiDock, assessing its effectiveness with large language models (LLMs) such as GPT-4 and analyzing the contribution of its components. Additionally, we compare FlakiDock with existing automated Dockerfile repair tools.
Our results show that FlakiDock achieves a 73.55% repair accuracy in resolving Dockerfile flakiness, significantly surpassing existing tools such as Parfum [4], which only achieves a 0.58% success rate. FlakiDock effectively addresses complex flakiness issues by employing static and dynamic information, similarity retrieval techniques, as well as an iterative feedback loop for refinement.
@@ II Motivation
Docker containerization aims to deliver a consistent and portable image build process. However, in practice, the Dockerfile build process often exhibits flakiness, posing a significant challenge to achieving these goals. This inconsistency can lead to several problems, such as unreliable deployments, increased debugging efforts, and unwanted delayed software delivery. Flakiness refers to the unpredictable nature of build failures in Dockerfiles over time, where a previously error-free Dockerfile may suddenly fail to build without any changes to the file or the project source code. This issue introduces challenges in maintaining reliable and efficient workflows. Despite the significance of this problem, there is a lack of comprehensive studies specifically addressing Dockerfile flakiness, which underscores the need for further research in this area.
As an illustrative example, Listing 1 represents a flaky Dockerfile, the corresponding build failure, and a subsequent repair based on the build error. According to the build output (lines 95 to 114), using a virtual environment is required to install pip for the specified alpine base image in Dockerfile (line 1) due to the adaptation of Python Enhancement Proposal (PEP) 668  [^2] https://peps.python.org/pep-0668/, which addresses Externally Managed Environments. This specification prevents package managers such as pip from modifying packages in the interpreter’s default environment, ensuring compatibility and reducing the risk of breaking the underlying operating system managed by external package managers. As a result of this adaptation, this Dockerfile fails to build, whereas previous builds prior to this adaptation were all successful. Thus, the solution here is to create and activate a virtual environment, as shown in the Dockerfile context (lines 11 and 12).
Existing Dockerfile static and dynamic analysis approaches assume the determinism of the Dockerfile behavior. Static analysis techniques [2, 9, 10, 4, 7] consider Dockerfile context to provide a set of smell patterns coherent with Docker best practices  [^3] https://docs.docker.com/build/building/best-practices/. The only Dynamic analysis approach [6], primarily focuses on the Dockerfile build output as an invariant to introduce and expand the error patterns. While adhering to best practices is essential to mitigate errors and vulnerabilities in Dockerfiles and in some cases to prevent potential failures from happening, we argue that this alone is insufficient to address flakiness. For example, existing tools such as HadoLint [9] recommend pinning the exact version of the base image (e.g., rule: DL3006) or dependencies (e.g., rules: DL3007, DL3008, DL3013, DL3016, DL3018) to prevent errors caused by their internal changes. This practice can be applied to Listing 1 by using old or outdated alpine images as a solution. However, it does not provide a viable solution for the problem; applying such rules without considering the static and dynamic nature of Dockerfiles can introduce other types of flakiness, such as outdatedness and compatibility issues in the future.
@@ 95 If the package in question is not packaged already (and
Dockerfiles rely on various elements such as operating systems, packages, environments, commands, and project source code, which can result in different forms of flakiness. Examples in Listing 1 and Listing 2 illustrate that understanding and resolving such flaky behavior requires analyzing both the static context (Dockerfile) and dynamic context (build output) along with its temporal changes. Currently, no existing study examines flakiness in Docker builds. To characterize the extent and causes of Docker build flakiness, we first conduct a longitudinal study of 18,055 Dockerized repositories over a period of nine months (section III). Second, we propose FlakiDock an automated approach for repairing Dockerfile flakiness utilizing both static and dynamic information (section IV).
@@ III Characterization of Dockerfile Flakiness
In this section, we delve into our empirical investigation of Dockerfile flakiness. We begin by detailing our data collection method and the setup for analyzing Dockerized projects for flaky behavior. Subsequently, we address two important research questions:
@@
  • RQ2: What are the categories of Dockerfile flakiness?
The entire analysis, including project checkouts and Dockerfile builds, is performed on an infrastructure consisting of 4 Intel(R) Xeon(R) CPU 2.50GHz machines with 62 GB RAM each.
@@ III-A Data Collection
For our study, we started with the Shipwright dataset [6], which contains 20,526 Docker projects with ten or more stars from GitHub repositories. The dataset includes Docker projects created up to June 2020 and focuses exclusively on projects with a single Dockerfile located in the root directory. This approach aims to facilitate analysis and mitigate challenges associated with multiple Dockerfiles within a repository.
Initially, we attempted to clone the most up-to-date version of all the repositories from the Shipwright dataset. However, some repositories were no longer publicly accessible, had been removed from GitHub, or no longer contained the root Dockerfiles. Consequently, our final dataset comprised 18,055 repositories that met the criteria and were available for analysis.
In our study, we analyze Dockerfiles along with their build outputs, which we generate by building the Dockerfiles within our infrastructure systems. Given that some Docker projects can be time-consuming to build, we set a 30-minute build timeout for each repository. This resulted in 93% of builds completing without a timeout. To ensure the reliability and efficiency of our large-scale Docker build system, we use the docker build command with the --no-cache option to eliminate failures stemming from cached Docker data. Additionally, we develop a systematic Docker cleaning technique to prevent environmental and internal errors, ensuring a fresh Docker environment. To enhance time efficiency and prevent internal errors, we clean the Docker system after every four consecutive builds—a frequency determined through trial and error. The cleaning process involves removing all the cache, images, and any other peripheral leftover data during the builds alongside uninstalling and reinstalling Docker with a steady version to prevent working with a corrupted Docker system. This pipeline ensures the freshness of the Docker system throughout the builds. Using this approach, we stored 32 rounds of builds for the initial Docker projects, capturing Dockerfile build outputs in our infrastructure machines from April 2023 to December 2023.
@@ III-B Flakiness Extraction
Filtering Phase. To uncover flakiness, we initiated an in-context Docker build for all the Dockerized GitHub projects in our dataset, of which (8,132 (45.04%)) were built successfully, and (9,923 (54.96%)) failed during the build procedure. We consider only the successful ones as valid candidates for our flakiness analysis due to their initial stability. Over nine months, we rebuilt the candidate projects 32 times (3.5 times per month on average). The reason for conducting builds over an extended timeframe was two-fold. First, building Dockerfiles is a time-consuming operation. The average time to build a single Dockerfile in our dataset is around eight minutes. Despite distributing our candidates across four different machines with similar operating systems and configurations and employing multiprocessing on each machine to enhance build speed, some degree of delay was unavoidable. Second, we expected that such an extended timespan could allow us to observe and investigate more fluctuations in project stability thoroughly.
Pre-Processsing Phase. The length of a Dockerfile build output varies based on the execution commands used, ranging from a few lines to tens of thousands, detailing the progress of each execution step. Therefore, diagnosing and extracting the important failure parts for further studies is essential. While extracting standard error logs can aid in understanding the root causes of issues, our extensive analysis revealed its inadequacy in capturing valuable insights in cases in which the error context is too long or the point of failure is far from the manifestation of error. To address such challenges, we implemented a rule-based pre-processing approach to extract error-related context from raw build outputs. In our method, we divide the build output into stages, each corresponding to an execution line in the Dockerfile. We examine each stage, capturing lines with error-related expressions alongside their adjacent lines with the same execution time to gather additional information about the captured errors.
@@ III-C Flakiness Categorization
Providing a taxonomy of Dockerfile flakiness serves as a vital step toward unraveling inconsistencies and complexities that come with flak Dockerfiles, thereby aiding in the identification and mitigation of the challenges developers face in practice. To achieve this, first, we use similarity-checking to remove duplicates or identical errors within each Dockerfile’s build outputs. Afterward, we perform an extensive study on pre-processed build outputs, leveraging LLMs to interpret and summarize the results. Furthermore, we utilize manual analysis to correct issues stemming from LLM misunderstandings.
Clustering Phase. During the rebuild period, we encountered 12,964 failing build outputs while building the candidate projects. To streamline our analysis and minimize redundancy within Dockerfile build outputs, we conduct a clustering process using sentence similarity assessment within each project. To identify unique build errors in a single project, we measure the cosine similarity between the build output embeddings, extracted from the sentence transformer: all-mpnet-base-v2  [^4] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 which is trained over 1 billion sentences in different domains. Using the similarity scores, we cluster them into distinct groups. In our clustering approach, each new build output is evaluated for its similarity to existing clusters. Based on the average similarity with all current clusters, we decide whether to incorporate it into an existing cluster or create a new one. We apply this clustering method within individual projects rather than across all projects due to the intricate nature of Dockerfiles and their build outputs, which complicates the accurate clustering of errors when applied globally. To enhance accuracy, we use pre-processed build outputs for similarity comparison rather than raw outputs. This method resulted in an 87% reduction in failing build outputs, leaving 1,684 remaining, which in turn simplifies subsequent analysis steps.

### v2

# Dockerfile Flakiness: Characterization and Repair

@@ Abstract
Dockerfile flakiness—unpredictable temporal build failures caused by external dependencies and evolving environments—undermines deployment reliability and increases debugging overhead. Unlike traditional Dockerfile issues, flakiness occurs without modifications to the Dockerfile itself, complicating its resolution. In this work, we present the first comprehensive study of Dockerfile flakiness, featuring a nine-month analysis of 8,132 Dockerized projects, revealing that around 10% exhibit flaky behavior. We propose a taxonomy categorizing common flakiness causes, including dependency errors and server connectivity issues. Existing tools fail to effectively address these challenges due to their reliance on pre-defined rules and limited generalizability. To overcome these limitations, we introduce FlakiDock, a novel repair framework combining static and dynamic analysis, similarity retrieval, and an iterative feedback loop powered by Large Language Models (LLMs). Our evaluation demonstrates that FlakiDock achieves a repair accuracy of 73.55%, significantly surpassing state-of-the-art tools and baselines.
@@
Docker, Flakiness, Large Language Models, Automated Program Repair
@@ I Introduction
Docker [1] is a set of platform-as-a-service (PaaS) products that use OS-level virtualization to automate building, deploying, and delivering applications in containers. This approach simplifies deployment across diverse systems and plays a critical role in the CI/CD pipeline, highlighting the importance of ensuring Dockerfile reliability during builds to maintain software dependability.
Existing academic [2, 3, 4, 5, 6, 7, 8, 9] and industrial efforts mainly focus on identifying Dockerfile smell or bug patterns. Static analysis studies [3, 10, 11, 8] offer a linter to detect smells inside Dockerfiles. While existing techniques effectively identify various Dockerfile smells, they often rely on pre-defined rules that are frequently outdated or poorly maintained [11]. Parfum [5] enriches the Dockerfile AST with structural information derived from command lines, enabling the automatic detection and repair of smells. Shipwright [7] employs a human-in-the-loop approach to repair broken Dockerfiles by clustering failure patterns using a modified BERT model [12] and HDBSCAN [13], then formalizing solutions into a repair pattern database with regular expressions and fix functions.
Unlike test flakiness [14], Dockerfile builds can exhibit temporal failures that emerge over time due to dynamic changes in external dependencies, such as updates to base images, third-party libraries, or environmental settings—occurring without any modifications to the Dockerfile itself. These temporal failures can manifest as either deterministic or non-deterministic behaviours, depending on the nature of the underlying issues. We refer to this phenomenon as Dockerfile flakiness—the unpredictable temporal behaviour of the Dockerfile build process, where builds may fail over time even though the content of the Dockerfile itself remains unchanged. Deterministic failures consistently cause builds to fail after a specific point, often due to changes such as the deprecation of a base image or the removal of a required dependency. In contrast, non-deterministic failures occur sporadically, resulting in random build successes and failures. Such issues are typically caused by transient factors, including server connectivity problems or temporary inconsistencies in package manager caches.
@@
Existing studies [2, 3, 4, 5, 6, 7, 8, 10, 11] assume the reliability of the Dockerfiles and their build output over time, so that a build failure indicates the changes applied directly to the source Dockerfile. The literature largely overlooks the issue of flakiness in Dockerfile builds, which can disrupt the deployment process, particularly in CI/CD pipelines, by hindering automatic builds. This leads to delays and consumes valuable developer time and effort in diagnosing and resolving the underlying causes of the flakiness.
@@
In this work, we first examine the consistency and reliability of Dockerfile builds over time. In a longitudinal study spanning a nine-month timeframe, we built and analyzed Dockerfiles of 8,132 open-source projects and observed that 798 (9.81%) exhibit flakiness behavior. Based on the observed instances of flakiness, we developed a taxonomy for characterizing Dockerfile flakiness. Our findings indicate that the most common types of flakiness are related to dependencies, web server connectivity, and security/authentication issues.
Second, we observe that existing techniques, such as Parfum and Shipwright, exhibit significant limitations in repairing flaky Dockerfiles. Parfum primarily addresses code smells, making it less effective in tackling flakiness issues. Shipwright, on the other hand, faces challenges including: (1) heavy reliance on human intervention, which hinders scalability and introduces variability due to dependence on the supervisor’s expertise, (2) limited generalizability of repair patterns to diverse or unseen errors, (3) reliance on pattern-matching and regular expressions, which can introduce inconsistencies, and (4) the need for ongoing maintenance of the repair database to accommodate emerging failure types.
To overcome these shortcomings, we present FlakiDock, a novel approach that leverages feedback-directed retrieval-augmented generation (RAG) with large language models (LLMs) to automatically repair Dockerfile flakiness. By integrating static and dynamic analysis, similarity retrieval, and an iterative LLM-based feedback loop, FlakiDock effectively resolves complex flakiness issues.
@@
  • The first study of Dockerfile flakiness characterization. We present the first taxonomy of Dockerfile flakiness by building Dockefiles and analyzing build outputs of 8,132 Dockerized projects.
  • A dataset called Flake4Dock including 798 and 7,325 flaky and non-flaky Dockerfiles. Flaky Dockerfiles are accompanied by categorization and build errors. We also provide repair information for 100 flaky Dockerfiles, specifically designed to evaluate flakiness detection and repair tasks.
  • An iterative LLM-based technique called FlakiDock that leverages both static and dynamic information from Dockerfiles to repair flakiness.
  • An empirical evaluation of FlakiDock, assessing its effectiveness with Dockerfile repair tools and baselines such as GPT-4.
Our results show that FlakiDock achieves a 73.55% repair accuracy in resolving Dockerfile flakiness, significantly outperforming Parfum (0.58%), Shipwright (5.52%), and GPT-4-based prompting with build output (37.79%).
@@ II Motivation
Listing 1 shows a flaky Dockerfile, the corresponding build failure, and a subsequent repair based on the build error. According to the build output (lines 95 to 114), using a virtual environment is required to install pip for the specified alpine base image in Dockerfile (line 1) due to the adaptation of Python Enhancement Proposal (PEP) 668 [15], which addresses Externally Managed Environments. This specification prevents package managers such as pip from modifying packages in the interpreter’s default environment, ensuring compatibility and reducing the risk of breaking the underlying operating system managed by external package managers. As such, this Dockerfile fails to build, whereas previous builds prior to this adaptation were all successful. The solution here would be to create and activate a virtual environment, as shown in the Dockerfile context (lines 11 and 12).
While adhering to best practices [3, 10, 11, 5, 8] is essential to mitigate errors and vulnerabilities in Dockerfiles and in some cases to prevent potential failures from happening, we argue that this alone is insufficient to address flakiness. For example, existing tools such as HadoLint [10] recommend pinning the exact version of the base image (e.g., rule: DL3006) or dependencies (e.g., rules: DL3007, DL3008, DL3013, DL3016, DL3018) to prevent errors caused by their internal changes. This practice can be applied to Listing 1 by using old or outdated alpine images as a solution. However, it does not provide a viable solution for the problem; applying such rules without considering the static and dynamic nature of Dockerfiles can introduce other types of flakiness, such as outdatedness and compatibility issues in the future.
@@ 95 If the package in question is not packaged already (and
Dockerfiles depend on various elements such as operating systems, packages, environments, commands, and project source code, making them susceptible to diverse forms of flakiness. Listings 1 and 2 demonstrate that understanding and resolving such flaky behavior requires analyzing both the static context (Dockerfile) and dynamic context (build output) along with its temporal changes.
@@ III Characterization of Dockerfile Flakiness
In this section we present our longitudinal investigation of Dockerfile builds, which addresses the following research questions:
@@
  • RQ2: What are the main categories of Dockerfile flakiness?
The entire analysis, including project checkouts and Dockerfile builds, was conducted on an infrastructure comprising four Intel Xeon 2.50GHz machines, each with 62 GB of RAM.
@@ III-A Data Collection
For our study, we started with the Shipwright dataset [7], which contains 20,526 Docker projects with ten or more stars from GitHub repositories. The dataset includes Docker projects created up to June 2020 and focuses exclusively on projects with a single Dockerfile located in the root directory. We cloned the most up-to-date version of all the repositories from the Shipwright dataset. However, some repositories were no longer publicly accessible, had been removed from GitHub, or no longer contained the root Dockerfiles. Consequently, our initial pool of projects comprised 18,055 repositories.
In our study, we analyze Dockerfiles along with their build outputs, which we generate by building the Dockerfiles within our infrastructure. Given that some Docker projects can be time-consuming to build, we set a 30-minute build timeout for each repository. This resulted in 93% of builds completing without a timeout. To ensure the reliability and efficiency of our large-scale Docker build system, we use the docker build command with the --no-cache option to eliminate failures stemming from cached Docker data. Additionally, we develop a systematic Docker cleaning technique to prevent environmental and internal errors, ensuring a fresh Docker environment. To enhance time efficiency and prevent internal errors, we clean the Docker system after every four consecutive builds—a frequency determined through trial and error. The cleaning process involves removing all the cache, images, and any other peripheral leftover data during the builds alongside uninstalling and reinstalling Docker with a steady version to prevent working with a corrupted Docker system. This pipeline ensures the freshness of the Docker system throughout the builds. Using this approach, we stored 32 rounds of builds for the initial Docker projects, capturing Dockerfile build outputs in our infrastructure machines from April 2023 to December 2023.
@@ III-B Flakiness Extraction
Filtering Phase. To start our flakiness study, we initiated an in-context Docker build for all the 18,055 Dockerized GitHub projects in our dataset that completed the build within 30-minutes. This resulted in 8,132 (45.04%) successful and 9,923 (54.96%) failed Docker builds. We consider only the successful ones as valid candidates for our flakiness analysis due to their initial stability.
Over a period of nine months, we rebuilt the 8,132 candidate projects weekly totaling 32 times. The reason for conducting builds over an extended timeframe was two-fold. First, building Dockerfiles is a time-consuming operation. The average time to build a single Dockerfile in our dataset is around eight minutes. Despite distributing our candidates across four different machines with similar operating systems and configurations and employing multiprocessing on each machine to enhance build speed, some degree of delay was unavoidable. Second, we expected that such an extended timespan could allow us to investigate temporal fluctuations in project stability.
Pre-Processsing Phase. Dockerfile build outputs can range from a few lines to thousands, detailing the progress of each execution step. Diagnosing and extracting critical failure points from these outputs is essential but challenging, especially when error logs are lengthy or the failure point is distant from the error’s manifestation. For instance, the RUN pip install -r requirements.txt command can produce hundreds of lines of output, often cluttered with warnings or unrelated messages, obscuring the root cause. To address this, we implemented a rule-based pre-processing approach that divides build outputs into stages, each corresponding to a specific Dockerfile command. Each stage includes lines containing the stage number, execution time, and log information. By analyzing each stage, we capture lines with error-related expressions like ParseError or ERR!, along with adjacent lines sharing the same execution time, to provide additional context for diagnosing failures.
@@ III-C Flakiness Categorization
A taxonomy of Dockerfile flakiness is critical for systematically addressing inconsistencies and complexities, enabling developers to identify and resolve related challenges. We employ similarity analysis to remove duplicates or identical errors within Dockerfile build outputs. This is followed by a detailed examination of pre-processed outputs, leveraging GPT-3.5 to interpret and summarize the results. Finally, manual analysis is performed to address potential inaccuracies.
Clustering Phase. During the rebuild period, we encountered 12,964 failing build outputs while building the candidate projects. To streamline our analysis and minimize redundancy within Dockerfile build outputs, we conduct a clustering process using sentence similarity assessment within each project. To identify unique build errors in a single project, we measure the cosine similarity between the build output embeddings, extracted from the all-mpnet-base-v2 [18] sentence transformer, which is trained over 1 billion sentences in different domains. Using the similarity scores, we cluster them into distinct groups. In our clustering approach, each new build output is evaluated for its similarity to existing clusters. Based on the average similarity with all current clusters, we decide whether to incorporate it into an existing cluster or create a new one. We apply this clustering method within individual projects rather than across all projects due to the intricate nature of Dockerfiles and their build outputs, which complicates the accurate clustering of errors when applied globally. To improve accuracy, we use pre-processed build outputs for similarity comparison rather than raw outputs. This approach resulted in an 87% reduction in failing build outputs, leaving 1,684 remaining, simplifying subsequent analysis steps.